OpenStudy (anonymous):

cos(π/∞)=0 ...Can anyone of you explain why cos(π/∞) is equal to 0 ? .......please give the solution also?

OpenStudy (anonymous):

there is no such number as \(\frac{\pi}{\infty}\)

OpenStudy (anonymous):

Like satellite said, that expression is not a number, but you probably mean to ask why \[\lim_{x\to\infty}\cos\frac{\pi}{x}=0\] which, by the way, is not a true statement. Consider the fraction \(\dfrac{\pi}{x}\). As \(x\to\infty\), the denominator gets larger and larger, so \(\dfrac{\pi}{x}\to0\). Since the cosine function is continuous, we can use the property that \[\lim_{x\to0}\cos(f(x))=\cos\left(\lim_{x\to0}f(x)\right)\] Since \(\dfrac{\pi}{x}\to0\), we're left with \[\cos\left(\lim_{x\to0}\frac{\pi}{x}\right)=\cos0=1\not=0\]